OpenStudy (anonymous):

5,8,11,14,17 write down an expression in terms of n for the nth term of this sequence

OpenStudy (anonymous):

t=5+(n-1)3 t=3n+2

OpenStudy (anonymous):

from the original formula for progressions t=a+(n-1)d

OpenStudy (saifoo.khan):

\[\large T_n = a(n-1)d\] where "a" is the first term. "n" is the nth term. "d" is the difference.